# Taming Multimodal Joint Training for High-Quality Video-to-Audio Synthesis

### v1

@@ Abstract
We propose to synthesize high-quality and synchronized audio, given video and optional text conditions, using a novel multimodal joint training framework (MMAudio). In contrast to single-modality training conditioned on (limited) video data only, MMAudio is jointly trained with larger-scale, readily available text-audio data to learn to generate semantically aligned high-quality audio samples. Additionally, we improve audio-visual synchrony with a conditional synchronization module that aligns video conditions with audio latents at the frame level. Trained with a flow matching objective, MMAudio achieves new video-to-audio state-of-the-art among public models in terms of audio quality, semantic alignment, and audio-visual synchronization, while having a low inference time (1.23s to generate an 8s clip) and just 157M parameters. MMAudio also achieves surprisingly competitive performance in text-to-audio generation, showing that joint training does not hinder single-modality performance. Code and demo are available at: hkchengrex.github.io/MMAudio.
@@ 1. Introduction
Current state-of-the-art video-to-audio methods either train only on audio-visual data [1] from scratch [63] or train new “control modules” [69, 18, 62, 65] for pretrained text-to-audio models on audio-visual data. The former is limited by the amount of available training data: the most commonly used dataset VGGSound [1] contains only around 550 hours of videos. Audio-visual data are expensive to curate at a large scale, as in-the-wild videos from the Internet 1) mostly contain music and speech111In AudioSet [8] (collected from YouTube), 82.48% of videos contain either human speech or music according to the provided audio labels., which has limited utility for training a general Foley model, and 2) contain non-diegetic [54] sounds such as background music or sound effects added in post-processing that are not suited for a Foley model. The latter, i.e., finetuning pretrained text-to-audio models (with some new adapter layers) on audio-visual data, enables to benefit from audio-generation knowledge obtained from larger-scale audio-text data (e.g., WavCaps [43] at 7,600 hours). However, adding control modules to pretrained text-to-audio models complicates the network architecture and limits the design space. It is also unclear whether pretrained text-to-audio models have sufficient degrees of freedom to support all video-to-audio scenarios, compared to training from scratch.
@@
To avoid these limitations, we propose a multimodal joint training paradigm (Figure 1) that jointly considers video, audio, and text in a single transformer network and masks away missing modalities during training. This allows us to train from scratch on both audio-visual and audio-text datasets with a simple end-to-end framework. Jointly training on large multimodal datasets enables a unified semantic space and exposes the model to more data for learning the distribution of natural audio. Empirically, with joint training, we observe a significant relative improvement in audio quality (10% lower Fréchet Distance [21] and 15% higher Inception Score [52]), semantic alignment (4% higher ImageBind [10] score), and temporal alignment (14% better synchronization score).
@@
To further improve temporal alignment, we note that using attention layers only is ineffective at indicating precise timings in audio/visual signals. To tackle this issue, instead, we introduce a conditional synchronization module that uses high frame-rate visual features (extracted from a self-supervised audio-visual desynchronization detector [17]) and operates in the space of scales and biases of adaptive layer normalization (adaLN) layers [46], leading to accurate frame-level synchronization (50% relative improvement in synchronization score).
@@
In summary, we first propose MMAudio, a multimodal joint training paradigm for video-to-audio. It enables accessible data scaling and cross-modal understanding, significantly improving audio quality and semantic alignment. We also propose a conditional synchronization module which enables more precise audio-visual synchrony. We train MMAudio on openly accessible datasets and scale it to two audio sampling rates (16kHz and 44.1kHz) and three model sizes (157M, 621M, 1.03B), with the smallest model already achieving new state-of-the-art performance in video-to-audio synthesis among public models. Surprisingly, our multimodal approach also achieves competitive performance in text-to-audio generation compared to dedicated text-to-audio methods, showing that joint training does not hinder single-modality performance. Our code and models will be released publicly.
@@ 2. Related Works
Semantic alignment between audio and video is obtained by training on paired audio-visual data with a generation objective [16, 5, 42, 62, 45, 56, 2, 67] or a contrastive objective [40]. To further understand the semantics of audio, we additionally train on paired audio-text data. We argue that the semantic understanding learned from audio-text pairs can be transferred to video-text pairs, as joint training leads to a shared semantic space (similar to ImageBind [10] and LanguageBind [70]) and enables the network to learn richer semantics from more diverse data.
@@
Besides learning temporal alignment directly from audio-visual pairs, some recent works first learn from videos a separate model to predict handcrafted proxy features such as onsets [69, 50], energy [18, 15], and root-mean-square (RMS) of waveforms [32, 3]. We deviate from these explicit handcrafted proxy features and directly learn alignment from the deep feature embeddings of a pre-trained self-supervised de-synchronization detector Synchformer [17], allowing a more nuanced interpretation of the input signal. A concurrent work V-AURA [61] also uses Synchformer [17] for synchronization in an autoregressive framework. However, V-AURA [61] does not perform multimodal training on text and has a short context window (2.56s) while we produce longer-term (8-10s) temporally consistent generation.
@@
The most common way to support multimodal conditioned generation is to add “control modules” to a pretrained text-to-audio network that injects visual and temporal features into the pretrained network [69, 18, 50, 32, 44, 15]. However, this usually further increases the number of parameters (the base text-to-audio network plus the control modules). Besides, as the text modality is fixed during video-to-audio training, it becomes more challenging to learn a joint semantic space – the video modality needs to bind to the semantics of text instead of both modalities learning to cooperate. In contrast, we train all modalities simultaneously in our multimodal training formulation to learn joint semantics and enable omnidirectional feature sharing among modalities. Alternatively, to align different modalities without training, Seeing-and-Hearing [65] uses a pretrained text-to-audio model and performs gradient ascent on an alignment score (e.g., ImageBind [10]) at test-time. We note that test time optimization is typically slower and sometimes results in low-quality and temporally misaligned output. Indeed, our model is faster at test time and is more consistent in producing synchronized audio. A concurrent work VATT [38] uses both video and text to generate audio and trains a captioning model to label unlabeled videos. However, they always require video condition inputs during training which limits the amount of training data. In contrast, our multimodal formulation allows us to train on audio-text datasets.
@@
Related to multimodal conditioning, multimodal generation models generate samples composed of multiple modalities (e.g., video and audio at the same time). We note that multimodal generation is a more challenging problem and existing approaches [23, 51, 58, 57] are not yet competitive with dedicated video-to-audio models. In this work, we focus on single-modal audio generation with multimodal conditions. We believe our multimodal formulation and architecture can serve as a foundation for future works in multimodal generation.
@@ 3.1. Preliminaries
For computational efficiency, we model the generative process in a latent space, following the common practice [36, 63]. For this, we first transform audio waveforms with the short-time Fourier transform (STFT) and extract the magnitude component as mel spectrograms [53]. The spectrograms are then encoded by a pretrained variational autoencoder (VAE) [25] into latents . During testing, the generated latents are decoded by the VAE into spectrograms which are then vocoded by a pretrained vocoder [33] into audio waveforms.
@@ 3.2. Overview
To estimate a flow for the current latent , MMAudio operates on the video/text conditions and the flow timestep . Figure 2 illustrates our network architecture. To combine inputs from different modalities, MMAudio consists of a series of () multimodal transformer blocks [7] with visual/text/audio branches, followed by a series of () audio-only transformer blocks [30]. Additionally, for audio-visual synchrony, we devise a conditional synchronization module that extracts and integrates into the generation process a high frame rate (24 frames per second (fps)) visual feature for temporal alignment. Next, we describe both components in detail.
@@ 3.3. Multimodal Transformer
Core to our approach is the desire to model the interactions between video, audio, and text modalities. For this purpose, we largely adopt the MM-DiT block design from SD3 [7] and introduce two new components for temporal alignment: aligned RoPE positional embeddings for aligning sequences of different frame rates and 1D convolutional MLPs (ConvMLPs) for capturing local temporal structure. Figure 2 (right) illustrates our block design. Note, we also include a sequence of audio-only single-modality blocks following FLUX [30], implemented by simply removing the streams of the two other modalities (joint attention becomes self-attention). Compared to considering all modalities at every layer, this design allows us to build a deeper network with the same parameter count and compute without sacrificing multimodality. This multimodal architecture allows the model to selectively attend to and focus on different modalities depending on the inputs, thus enabling effective joint training on both audio-visual and audio-text data. Next, we describe the feature representation in our network and then the core components of our block design.
@@
We represent all the features as one-dimensional tokens. Note, we deliberately do not use any absolute position encoding which allows us to generalize to different durations at test time. Thus, we specify the sequences in terms of frame rates to determine the number of tokens for a given duration. The visual features (one token per frame, at 8 fps) and text features (77 tokens) are extracted from CLIP [49] as 1024d features. The audio latents are in the VAE latent space (Section 3.1), at 31.25 fps as 20d latents by default. The synchronization features are extracted from Synchformer [17] at 24 fps as 768d features, which we will detail in Section 3.4. Note, except for the text tokens, all other tokens follow the same temporal ordering, albeit at different frame rates. After the initial ConvMLP/MLP layers, all features will be projected to the hidden dimension .
@@
These tokens from different modalities communicate via joint attention (Figure 2, right). Following [7], we concatenate the query, key, and value representations from the three different modalities and apply a joint attention operation [60]. The output is split into three modalities, following the input partition. We refer readers to [7] for details. We note that joint attention alone does not capture temporal alignment which we will address next.
@@
For audio-visual synchrony, precise temporal alignment is crucial. As typical in transformers [60], we adopt positional embeddings to inform the attention layers of time. Specifically, we apply the RoPE [55] embeddings on the queries and keys in the visual and audio streams before joint attention (Figure 2). Note that we do not apply it to the text stream since it does not follow the temporal order of video or audio. Further, since the frame rates do not align (8 fps for the visual stream, 31.25 fps for the audio stream), we scale the frequencies of the positional embeddings in the visual stream proportionally, i.e., by . We visualize the default (non-aligned) RoPE and our proposed aligned RoPE in Figure A5. We note that these aligned embeddings are beneficial yet insufficient for good synchrony. Therefore, we introduce an additional synchronization module, which we discuss in Section 3.4.
@@ 3.4. Conditional Synchronization Module
The upsample operation uses nearest neighbor interpolation and matches the frame rate of the synchronization features with that of the audio latent . This frame-aligned conditioning is injected via the adaLN layers in the audio stream for precise feature modulation. Similar to Eq. 4, we apply via
@@ 3.5.1. Multimodal Datasets
We train on VGGSound [1] as the only audio-visual dataset. It offers around 500 hours of footage. Additionally, VGGSound contains a class label (310 classes in total) for each video and we use the class names as input following ReWaS [18] and VATT [38]. We set aside 2K videos from the training set for validation, resulting in a training set of around 180K 10s videos. We truncate the videos to 8s from the end for training.
@@
We use AudioCaps [22] (around 128 hours, manually captioned) and WavCaps [43] (around 7,600 hours, automatically captioned from metadata) as audio-text datasets for training. Since they do not contain the visual modality, we set all the visual features and synchronization features corresponding to these samples as learnable empty tokens and respectively. For short audios (<16s), we truncate them to 8s from the end for training as in VGGSound. For longer audios, we take at most five non-overlapping crops of 8s from each audio. This results in a total of 951K audio clip-text pairs.
@@
We notice a minor (<1% of the test sets) train/test data contamination among these datasets. For a fair comparison, we have removed the test sets of VGGSound and AudioCaps from all training data. We provide more details in Appendix C.
@@ 3.5.2. Implementation Details
Our default model generates 16kHz audio encoded as 20-dimensional, 31.25fps latents (following Frieren [63]), with . We refer to this default model as ‘S-16kHz’. We additionally train larger models and models with higher audio sampling rates: ‘S-44.1kHz’, ‘M-44.1kHz’, and ‘L-44.1kHz’, detailed in Section E.1. The parameter counts and running time of these models are summarized in Section 4. We describe additional implementation details in Appendices F, D and E.
@@
To enable classifier-free guidance [12] in inference, we randomly mask away all visual tokens ( and ) or all the text with a probability during training. The masked visual tokens are replaced with a learnable embedding ( and ), while any masked text is replaced with the empty string .
@@ 4.2. Main Results
Section 4 compares our main results on the VGGSound [1] test set (15K videos) with existing state-of-the-art models. We evaluate all generations at 8s following Wang et al. [63] by truncating longer audio to 8 seconds. For V-AURA [61], we use the official autoregression code to generate 8s audio. ReWaS [18] only generates 5s audio thus we evaluate it as-is, by truncating the ground truth also to 5s – we indicate this discrepancy via the gray font in the table. Our smallest model (157M) demonstrates better distribution matching, audio quality, semantic alignment, and temporal alignment than prior methods, while being fast. A notable exception is the IB-score comparison with Seeing-and-Hearing [65]. We note that Seeing-and-Hearing [65] directly optimizes the IB score during test time, which we do not perform. Further, our larger models continue to improve in FD and IB-score, though we observe diminishing returns potentially limited by data quality and the amount of audio-visual data. Note, while our method uses more data for multimodal joint training, we do not use more data overall than some of the existing methods: FoleyCrafter [69], V2A-Mapper [62], ReWaS [18], and Seeing-and-Hearing [65] all finetune/incorporate a text-to-audio model that has been trained on audio-text data similar to the one we use. For a fair evaluation, we use the precomputed samples provided by [63, 62, 38], and reproduce the results using the official inference code for [61, 69, 18, 65]. Figure 3 visualizes our results and compares them with prior works. We present the results of a user study and comparison with Movie Gen Audio [48] in Appendices A and B.

### v2

# MMAudio: Taming Multimodal Joint Training for High-Quality Video-to-Audio Synthesis

@@ Abstract
We propose to synthesize high-quality and synchronized audio, given video and optional text conditions, using a novel multimodal joint training framework (MMAudio). In contrast to single-modality training conditioned on (limited) video data only, MMAudio is jointly trained with larger-scale, readily available text-audio data to learn to generate semantically aligned high-quality audio samples. Additionally, we improve audio-visual synchrony with a conditional synchronization module that aligns video conditions with audio latents at the frame level. Trained with a flow matching objective, MMAudio achieves new video-to-audio state-of-the-art among public models in terms of audio quality, semantic alignment, and audio-visual synchronization, while having a low inference time (1.23s to generate an 8s clip) and just 157M parameters. MMAudio also achieves surprisingly competitive performance in text-to-audio generation, showing that joint training does not hinder single-modality performance. Code, models, and demo are available at: hkchengrex.github.io/MMAudio.
@@ 1. Introduction
Current state-of-the-art video-to-audio methods either train only on audio-visual data [1] from scratch [67] or train new “control modules” [73, 20, 66, 69] for pretrained text-to-audio models on audio-visual data. The former is limited by the amount of available training data: the most commonly used audio-visual dataset VGGSound [1] contains only around 550 hours of videos. Audio-visual data are expensive to curate at a large scale, as in-the-wild videos from the Internet 1) contain music and speech111In AudioSet [9] (collected from YouTube), 82.48% of videos contain either human speech or music according to the provided audio labels., which has limited utility for training a general Foley model, and 2) contain non-diegetic [58] sounds such as background music or sound effects added in post-processing, again unsuitable for a Foley model. The latter, i.e., finetuning pretrained text-to-audio models (with added control modules) on audio-visual data, enables models to benefit from audio-generation knowledge obtained from larger-scale audio-text data (e.g., WavCaps [46] at 7,600 hours). However, adding control modules to pretrained text-to-audio models complicates the network architecture and limits the design space. It is also unclear whether pretrained text-to-audio models have sufficient degrees of freedom to support all video-to-audio scenarios, compared to training from scratch.
@@
To avoid these limitations, we propose a multimodal joint training paradigm (Figure 1) that jointly considers video, audio, and text in a single multimodal transformer network and masks missing modalities during training. This allows us to train from scratch on both audio-visual and audio-text datasets with a simple end-to-end framework. Jointly training on large multimodal datasets enables a unified semantic space and exposes the model to more data for learning the distribution of natural audio. Empirically, with joint training, we observe a significant relative improvement in audio quality (10% lower Fréchet Distance [23] and 15% higher Inception Score [56]), semantic alignment (4% higher ImageBind [11] score), and temporal alignment (14% better synchronization score).
@@
To further improve temporal alignment, we introduce a conditional synchronization module that uses high frame-rate visual features (extracted from a self-supervised audio-visual desynchronization detector [19]) and operates in the space of scales and biases of adaptive layer normalization (adaLN) layers [50], leading to accurate synchronization (50% relative improvement in synchronization score).
@@
In summary, we first propose MMAudio, a multimodal joint training paradigm for video-to-audio. It enables accessible data scaling and cross-modal understanding, significantly improving audio quality and semantic alignment. We also propose a conditional synchronization module which enables more precise audio-visual synchrony. We train MMAudio on openly accessible datasets and scale it to two audio sampling rates (16kHz and 44.1kHz) and three model sizes (157M, 621M, 1.03B), with the smallest model already achieving new state-of-the-art performance in video-to-audio synthesis among public models. Surprisingly, our multimodal approach also achieves competitive performance in text-to-audio generation compared to dedicated text-to-audio methods, showing that joint training does not hinder single-modality performance.
@@ 2. Related Works
Semantic alignment between audio and video is learned by training on paired audio-visual data with a generation objective [18, 6, 44, 66, 49, 60, 2, 71] or a contrastive objective [42]. To further understand audio semantics, we additionally train on paired audio-text data. We argue that the semantic understanding learned from audio-text pairs can be transferred to video-text pairs, as joint training leads to a shared semantic space (similar to ImageBind [11] and LanguageBind [74]) and enables the network to learn richer semantics from more diverse data.
@@
Besides learning temporal alignment directly from audio-visual pairs, some recent works first learn from videos a separate model to predict handcrafted proxy features such as audio onsets [73, 54], energy [20, 17], or root-mean-square of waveforms [34, 4]. We deviate from these handcrafted features and directly learn alignment from the deep feature embeddings of a pre-trained self-supervised desynchronization detector Synchformer [19], allowing a more nuanced interpretation of the input signal. A recent work V-AURA [65] also uses Synchformer [19] for synchronization in an autoregressive framework. However, [65] does not perform multimodal training on text and has a short context window (2.56s) while we produce longer-term (8-10s) temporally consistent generations. In terms of improving positional embeddings for temporal alignment, Mei et al. [45] concurrently propose to subsample high-frequency (audio) positional embeddings, while we propose to scale up the frequencies of low-frequency (visual) positional embeddings – the effects are identical when the higher frequency is an integer multiple of the lower one.
@@
The most common way to support multimodal conditioned generation is to add “control modules” that inject visual features to a pretrained text-to-audio network [73, 20, 54, 34, 47, 17, 13]. However, this increases the number of parameters. Besides, as the text modality is fixed during video-to-audio training, it becomes more challenging to learn a joint semantic space – the video modality needs to bind to the semantics of text instead of both modalities learning to cooperate. In contrast, we train all modalities simultaneously in our multimodal training formulation to learn joint semantics and enable omnidirectional feature sharing among modalities. Alternatively, to align different modalities without training, Seeing-and-Hearing [69] uses a pretrained text-to-audio model and performs gradient ascent on an alignment score (i.e., ImageBind [11]) at test-time. We note that the test time optimization is slow and sometimes results in low-quality and temporally misaligned output. Indeed, our model is faster at test time and more consistently produces synchronized audio. Concurrent to our work, VATT [40] and MultiFoley [3] explore jointly trained multimodal conditioning. VATT [40] uses both video and text to generate audio but always requires video conditions during training. MultiFoley [3] is formulated similarly to ours, but MMAudio uses much higher frame-rate visual features (24 FPS, while MultiFoley uses 8 FPS features), which leads to significantly better audio-visual synchronization.
@@
Related to multimodal conditioning, multimodal generation models produce samples composed of multiple modalities (e.g., video and audio). Multimodal generation is more challenging and existing approaches [25, 55, 62, 61] are not yet competitive with dedicated video-to-audio models. In this work, we focus on multimodal conditioned audio generation. We believe our multimodal formulation and architecture serve as a foundation for future work in multimodal generation.
@@ 3.1. Preliminaries
For computational efficiency, we model the generative process in a latent space, following the common practice [38, 67]. For this, we first transform audio waveforms via Short-Time Fourier Transform (STFT) and extract mel spectrograms [57], which are then encoded by a pretrained variational autoencoder (VAE) [27] into latents . During testing, the generated latents are decoded by the VAE into spectrograms which are then vocoded by a pretrained vocoder [35] into audio waveforms.
@@ 3.2. Overview
To estimate a flow for the current latent , MMAudio operates on the video/text conditions and the flow timestep . Figure 2 illustrates our network architecture. To combine inputs from different modalities, MMAudio consists of a series of () multimodal transformer blocks [8] with visual/text/audio branches, followed by a series of () audio-only transformer blocks [32]. Additionally, for audio-visual synchrony, we devise a conditional synchronization module that extracts and integrates into the generation process high frame rate (24 frames per second (fps)) visual features for temporal alignment. Next, we describe both components in detail.
@@ 3.3. Multimodal Transformer
Core to our approach is the desire to model the interactions between video, audio, and text modalities. For this purpose, we largely adopt the MM-DiT block design from SD3 [8] and introduce two new components for temporal alignment: aligned RoPE positional embeddings for aligning sequences of different frame rates and 1D convolutional MLPs (ConvMLPs) for capturing local temporal structure. Figure 2 (right) illustrates our block design. Note, we also include a sequence of audio-only single-modality blocks following FLUX [32], implemented by simply removing the streams of the two other modalities (i.e., the joint attention becomes a self-attention). Compared to considering all modalities at every layer, this design allows us to build a deeper network with the same parameter count and compute without sacrificing multimodality. This multimodal architecture allows the model to selectively attend to and focus on different modalities depending on the inputs, thus enabling effective joint training on both audio-visual and audio-text data. Next, we describe the feature representation in our network and then the core components of our block design.
@@
We represent all the features as one-dimensional tokens. Note, we deliberately do not use any absolute position encoding which allows us to generalize to different durations at test time. Thus, we specify the temporal sequences in terms of frame rates to determine the number of tokens for a given duration. The visual features (one token per frame, at 8 fps) and text features (77 tokens) are extracted from CLIP [53] as 1024d features. The audio latents are in the VAE latent space (Section 3.1), at 31.25 fps as 20d latents by default. The synchronization features are extracted with Synchformer [19] at 24 fps as 768d features, which we will detail in Section 3.4. Note, except for the text tokens, all other tokens follow the same temporal ordering, albeit at different frame rates. After the initial ConvMLP/MLP layers, all features will be projected to the hidden dimension .
@@
These tokens from different modalities communicate via joint attention (Figure 2, right). Following [8], we concatenate the query, key, and value representations from the three different modalities and apply scaled dot product attention [64]. The output is split into three modalities, following the input partition. We refer readers to [8] for details. We note that joint attention alone does not capture temporal alignment which we will address next.
@@
For audio-visual synchrony, precise temporal alignment is crucial. As typical in transformers [64], we adopt positional embeddings to inform the attention layers of time. Specifically, we apply RoPE [59] embeddings on the queries and keys in the visual and audio streams before joint attention (Figure 2). Note that we do not apply it to the text stream since it does not follow the temporal order of video or audio. Further, since the frame rates do not align (8 fps for the visual stream, 31.25 fps for the audio stream), we scale the frequencies of the positional embeddings in the visual stream proportionally, i.e., by . We visualize the default (non-aligned) RoPE and our proposed aligned RoPE in Figure A5. We note that these aligned embeddings are beneficial yet insufficient for good synchrony. Therefore, we introduce an additional synchronization module, which we discuss in Section 3.4.
@@ 3.4. Conditional Synchronization Module
Upsample uses nearest neighbor interpolation and matches the frame rate of the synchronization features with that of the audio latent . This frame-aligned conditioning is injected via the adaLN layers in the audio stream for feature modulation. Similar to Eq. 4, we apply via
@@ 3.5.1. Multimodal Datasets
We train on VGGSound [1] as the only audio-text-visual dataset. It offers around 500 hours of footage. Additionally, VGGSound contains a class label (310 classes in total) for each video and we use the class names as input following ReWaS [20] and VATT [40]. We set aside 2K videos from the training set for validation, resulting in a training set of around 180K 10s videos. We use the first 8s of each video for training.
@@
We use AudioCaps [24] (128 hours, manually captioned), Clotho [5] (31 hours, manually captioned), and WavCaps [46] (7,600 hours, automatically captioned from metadata) as audio-text datasets for training. Since they do not contain the visual modality, we set all the visual features and synchronization features corresponding to these samples as learnable empty tokens and respectively. For short audios (<16s), we truncate them to 8s for training, as in VGGSound. For longer audios, we take up to five non-overlapping crops of 8s each. This results in a total of 951K audio clip-text pairs.
@@
We notice a minor (<1% of the test sets) train/test data contamination among these datasets. For a fair comparison, we have removed the test sets of VGGSound and AudioCaps from all training data. We provide more details in Appendix E.
@@ 3.5.2. Implementation Details
Our default model generates 16kHz audio encoded as 20-dimensional, 31.25fps latents (following Frieren [67]), with . We refer to this default model as ‘S-16kHz’. We additionally train larger models and models with higher audio sampling rates: ‘S-44.1kHz’, ‘M-44.1kHz’, and ‘L-44.1kHz’, detailed in Section G.1. The parameter counts and running time of these models are summarized in Section 4. We describe additional implementation details in Appendices H, F and G.
@@
To enable classifier-free guidance [14] during inference, we randomly mask away the visual tokens ( and ) or the text with a probability during training. The masked visual tokens are replaced with learnable tokens ( and ), while any masked text is replaced with the empty string .
@@ 4.2. Main Results
Section 4 compares our main results on the VGGSound [1] test set (15K videos) with existing state-of-the-art models. We evaluate all generations at 8s following Wang et al. [67] by truncating longer audio to 8 seconds. For V-AURA [65], we use the official autoregression code to generate 8s audio. ReWaS [20] only generates 5s audio thus we evaluate it as-is, by truncating the ground truth also to 5s – we indicate this discrepancy via the gray font in the table. Our smallest model (157M) demonstrates better distribution matching, audio quality, semantic alignment, and temporal alignment than prior methods, while being fast. A notable exception is the IB-score comparison with Seeing-and-Hearing [69] (SAH). We note that SAH directly optimizes the IB score during test time, which we do not perform. Further, our larger models continue to improve in FD and IB-score, though we observe diminishing returns potentially limited by data quality and the amount of audio-visual data. Note, while our method uses more data for multimodal joint training, we do not use more data overall than some of the existing methods: FoleyCrafter [73], V2A-Mapper [66], ReWaS [20], and SAH [69] all finetune/incorporate a text-to-audio model that has been trained on audio-text data similar to the one we use. For a fair evaluation, we use the precomputed samples provided by [67, 66, 40], and reproduce the results using the official inference code for [65, 73, 20, 69]. Figure 3 visualizes our results and compares them with prior works. We present the results of a user study and comparison with Movie Gen Audio [52] in Appendices A and B. To address any potential bias introduced by using the model-based DeSync metric, we additionally assess synchronization via model-free metrics on Greatest Hits [48] in Section 4.2, detailed in Appendix C.
l@ c@ c@ c@ c
Method Acc. AP F1 DeSync
Frieren [67] 0.6949 0.7846 0.6550 0.851
V-AURA [65] 0.5852 0.8567 0.6441 0.654
FoleyCrafter [73] 0.4533 0.6939 0.4319 1.225
Seeing&Hearing [69] 0.1156 0.8342 0.1591 1.204
MMAudio-S-16kHz 0.7637 0.9010 0.7928 0.483
MMAudio-S-44.1kHz 0.7150 0.9097 0.7666 0.444
MMAudio-M-44.1kHz 0.7226 0.9054 0.7620 0.443
MMAudio-L-44.1kHz 0.7158 0.9064 0.7535 0.442